# Detecting Offensive Memes with Social Biases in Singapore Context Using Multimodal Large Language Models

### v1

@@ Abstract.
The dataset, code and model weights will be open-sourced at https://github.com/aliencaocao/vlm-for-memes-aisg.
@@ 1.2. Contributions
We gather, filter and annotate an instruction fine-tuning dataset with three splits, totaling 112277 multimodal training samples (global-context memes and Singapore-context memes) and 715 Singapore-related, mixed-modality Wikipedia corpus. The dataset aims to teach VLMs to classify memes that are offensive and unsafe for social media. Each sample contains an image of the meme, textual descriptions for VLM reasoning, and a binary label stating if the meme is offensive or not. The dataset and any processing code associated with them will be open sourced.
@@ 4.1. Overview
We list each component of our combined dataset in Table 1. Unless otherwise specified, all datasets have either an unspecified license, or are considered to be in the public domain. Datasets beginning with ”@” were scraped from public Instagram pages using Instaloader(Graf, 2024). Datasets containing Singapore-context memes are labeled with ”(SG)”. We remove duplicates within each dataset using image filename. A total of 2897 samples were taken from the pool of Singapore-context memes as a held-out validation set, which leaves a total training dataset of 109380 memes.
@@ 4.4. Singapore-specific meme datasets
To provide the Singapore context and up-to-date memes, we scraped internet platforms such as the r/Singapore subreddit, Facebook and Instagram accounts posting various themes for memes. Table 2 summarizes these accounts. The scraped r/Singapore dataset contains 1461 samples that are posted with the ”Meme” and ”SHITPOST” flair, after removing duplicates and deleted posts. As there were no labels for scraped data, all labels were fully generated by GPT-4V. We carefully chose these popular-among-locals sources focusing on different themes to ensure that memes collected from them boasts high diversity while being localized and aligned with modern societal interests.
@@ 5.1. OCR
We employ optical character recognition (OCR) as a first step to improve the model’s ability to read the text in the sample. This is achieved using the PaddleOCR library, which offers fast multilingual character detection. A first pass is done with the multilingual detection model PP-OCRv4 (successor of PP-OCRv3(Li et al., 2022) to detect characters. The bounding boxes are passed to PaddleClas’s PPLCNet_x1_0(Cui et al., 2021), which detects the language script (Latin, Chinese, or Tamil). The appropriate language-specific OCR model is then used to provide the final results, to ensure highest accuracy. If any text is detected, abbreviation expansion will be done as mentioned in subsection 4.5.
@@ 8. Acknowledgments
Jiayang, Yuxuan, Sherman and Alistair contributed to the collection and processing of meme datasets. Theodore contributed to Wikipedia and Instagram scraping. Jiayang led training efforts for LLaVA models while Yuxuan led training efforts of Qwen2-VL models and evaluation of both VLMs. Jiayang, Yuxuan and Sherman contributed to the inference pipeline and its performance optimizations. Sherman and Bryan provided training-related advisories and compute resources.

### v2

@@ Abstract.
The dataset, code and model weights have been open-sourced at https://github.com/aliencaocao/vlm-for-memes-aisg.
@@ 1.2. Contributions
We gather, filter, and annotate an instruction fine-tuning dataset with three splits, totaling 112277 global and Singapore-context memes and 715 Singapore-related, mixed-modality Wikipedia corpus. The dataset aims to teach VLMs to classify memes that are offensive and unsafe for social media. Each sample contains an image of the meme, textual descriptions for VLM reasoning, and a binary label stating if the meme is offensive or not. The dataset and any processing code associated with them will be open sourced.
@@ 2. Related Work
2.1. Singlish in LLMs
There are some recent works focusing on fine-tuning LLMs to understand Singlish texts. AI Singapore has led the training efforts on SEA-LION (AISingapore, 2024), a series of models focusing on Southeast Asia languages through continued pre-training on Southeast Asian languages. However, their efforts are centralized around unimodal text understanding, and cannot be directly applied to meme classification without significant fine-tuning to re-align the vision features. GovTech Singapore developed LionGuard (Foo and Khoo, 2024) which specializes in moderating texts in Singlish and Singaporean context. Similar to SEA-LION, LionGuard is text-only and uses text embedding models and linear probing to perform the classification. This approach is more compute-efficient than token generation and could be applied to multimodal contents like memes. However, current multimodal embedding models are still insufficiently explored and evaluated. They also require significantly more data to fine-tune for localization, which is hard for a low-resource language. LionGuard also highlighted the importance of localization in content moderation, especially on slangs and words that are only offensive in the Singapore context. This aligns with our findings.
@@ 4.1. Overview
We list each component of our combined dataset in Table 1. Unless otherwise specified, all datasets have either an unspecified license, or are considered to be in the public domain. Datasets beginning with ”@” were scraped from public Instagram pages using Instaloader(Graf, 2024). Datasets containing Singapore-context memes are labeled with ”(SG)”, which are from sources that posts Singapore-context content only, and target Singaporeans as the audience. We remove duplicates within each dataset using image filename. A total of 2897 Singapore-context memes were taken out as a held-out validation set, which leaves a training set of 109380 memes.
@@ 4.4. Singapore-specific meme datasets
To provide the Singapore context and up-to-date memes, we scraped internet platforms such as the r/Singapore subreddit, Facebook and Instagram accounts posting various themes for memes. Table 2 summarizes these accounts. The scraped r/Singapore dataset contains memes that were posted with the ”Meme” and ”SHITPOST” flairs. As there were no labels for scraped data, all labels were fully generated by GPT-4V. We carefully chose these popular-among-locals sources focusing on different themes to ensure that memes collected from them boasts high diversity while being localized and aligned with modern societal interests.
@@ 5.1. OCR
We employ optical character recognition (OCR) as a first step to improve the model’s ability to read the text in the sample. This is achieved using the PaddleOCR library, which offers fast multilingual character detection. A first pass is done with the multilingual detection model PP-OCRv4 (successor of PP-OCRv3(Li et al., 2022)) to detect characters. The bounding boxes are passed to PaddleClas’s PPLCNet_x1_0(Cui et al., 2021), which detects the language script (Latin, Chinese, or Tamil). In case of multiple scripts detected, the majority will be chosen. If the detected script is not one of the three, it is treated as English and passed directly to VLM. We estimate that the frequency of this happening is extremely low and does not warrant additional compute resources for additional processing. The appropriate language-specific OCR model is then used to provide the final results, to ensure highest accuracy. If any text is detected, abbreviation expansion will be done as mentioned in subsection 4.5.
@@ 5.6. Evaluation
We evaluated our methods on the above-mentioned validation set consisting 2897 Singapore-context memes. We could not evaluate on any existing global-context meme datasets as they have all been used for training.
@@ 7.3. Future Work
Human expert-labeled test set could also prove valuable to evaluate our work with greater accuracy and reliability, as the current evaluation depends on the underlying MLLM used to label the dataset.
@@ 8. Acknowledgments
Jiayang, Yuxuan, Sherman and Alistair contributed to the collection and processing of meme datasets. Theodore contributed to Wikipedia and Instagram scraping. Jiayang led training efforts for LLaVA models while Yuxuan led training efforts of Qwen2-VL models and evaluation of both VLMs. Jiayang, Yuxuan and Sherman contributed to the inference pipeline and its performance optimizations. Sherman and Bryan provided training-related advisories and compute resources. Yuxuan, Bryan and Alistair contributed to writing the paper.